You are going on a dream vacation with five of your closest friends for a week. Since you are taking a math course, your friends asked you to come up with a budget to see how much it will cost each of you. Please identify your location in the US (input in cell D2) and then determine the total costs (for everyone going) remember to be thorough so the group does not run out of money. Your expense list must include the following (this means you cannot book an all expenses paid trip) travel expenses, such as plane or bus fare, local travel within the area. Lodging expenses. Food expenses. Entertainment or activity expenses.
1
Expert's answer
2020-05-06T16:54:41-0400


Let's say 5 friends fly from Boston to New York

1.Ticket price from Boston to New York: $131

"131\times 5=655"

2.Hotel accommodation for 7 days:

"380.17*7=2661.19"

3.Food in the city:

average check per person in one day (breakfast, lunch, dinner):

=8.95+9.95+4.25+7.95+7.95+7.95=47,

"47*5*7=1645"

4.Entertainment in the city:

First two days just walks in new york:

Brooklyn Bridge, Times Square, Wall Street, Soho District, New York Public Library, Central Station, Brooklyn Botanical Garden

Statue of Liberty:

"18\times5=90"

Metropolitan Museum:

"25\times5=125"

The Museum of Modern Art

"25\times5=125"

Museum of Memory 9/11

"24\times5=120"

Rockefeller Center

"52\times5=260"

5. New York Boston plane ticket:

"135*5=675"


Total costs: 655+2661.19+1645+90+125+125+120+260+675=6356.19

Total cost per person

"\frac{6356.19}{5}=1271.24"


The best time to visit New York is in autumn and spring. Have a nice trip!